OpenStudy (anonymous):

what is the factored binomial from the trinomial x^2+x-12?

OpenStudy (anonymous):

what can multiply to make 12 4 and 3 6 and 2 1 and 12 which one when you subtract would give you either 1 or -1?

OpenStudy (anonymous):

subtract each pair, for ex 4-3 6-2 1-12

OpenStudy (anonymous):

4-3 =1 which the same as the middle factor 1*x. let's try 4 and 3 as our factor (x+4)(x-3) Now check if it's correct (x+4)(x-3) =x^2 +4x -3x -12 = x^2 + x -12 Checked.

OpenStudy (anonymous):

thanks